OpenStudy (anonymous):

y=6(10a+4b) solve the following formula for a

OpenStudy (anonymous):

y=60a+24b y-24b = 60a \[\frac{ y-24b }{ 60 } = a\]

OpenStudy (anonymous):

thanks